OpenStudy (anonymous):

Write the equation of the quadratic function with roots -1 and -7 and a vertex at (-4, 7)

OpenStudy (anonymous):

Roots at -1 and -7 so will have the form: \[y=a(x+1)(x+7)\] Substitute x=-4, y=7 to find a. Can you have a go?

OpenStudy (anonymous):

\[y=a(x+1)(x+7)\] So that when you set y=0 (i.e. on the x-axis), x satisfies \[0=a(x+1)(x+7)\] when it is equal to -1 or -7

OpenStudy (anonymous):

7=a(-4+1)(-4+7)

OpenStudy (anonymous):

Yep, so: \[7=a\times(-3)\times(3)\] \[7=-9a\] \[a=\frac{-7}{9}\] Now do you see what the equation will be with this value of a?

OpenStudy (anonymous):

y=7(x+1)(x=7) ?

OpenStudy (anonymous):

@Traxter

OpenStudy (anonymous):

Not quite, if you substitue a=-7/9 into the equation I wrote before: \[y=a(x+1)(x+7)\] Then you will have: \[y=\frac{-7}{9}(x+1)(x+7)\] Agree?

OpenStudy (anonymous):

yes!!

OpenStudy (anonymous):

thank youuuuuuuu

OpenStudy (anonymous):

and @henpen

OpenStudy (anonymous):

Welcome :)